# Model Collapse in the Self-Consuming Chain of Diffusion Finetuning: A Novel Perspective from Quantitative Trait Modeling

### v2

@@ A.1 Hyperparameters
We finetune Stable Diffusion v1.5 [24] using LoRA [25] at each iteration, with ft-MSE [48] as a fixed VAE to project images into latent space. Horizontal flip is the only image augmentation applied. Our implementation follows [26] and is built on PyTorch v2.2.2 [49], with torchvision 0.17.2, running on CUDA 12.4 using NVIDIA A-100 and L40S GPUs. All default hyperparameters are listed in Table 1.
@@ A.2 Datasets
Butterfly [29] includes $832$ images across $8$ species, using BLIP captions (length $50$-$75$ words) and species descriptions.
Appendix B Different CFG Scales
This section presents images from Chain of Diffusion at various CFG scales for the four datasets. Figure 8, 9, 10, and 11 corresponds to Pokemon, CelebA-1k, Kumapi, and Butterfly, respectively. Each dataset shows results for five CFG scales, covering high, medium, and low values. The optimal medium CFG scale is $2.5$ for Pokemon and Kumapi, and $1.5$ for CelebA-1k and Butterfly. Lower-than-optimal CFG scales lead to low-frequency degradations, while higher-than-optimal scales result in high-frequency degradations with saturated colors and repetitive patterns. Notably, the optimal CFG for Pokemon and Kumapi causes severe degradation in CelebA-1k and Butterfly.
@@ Appendix B Different CFG Scales
Appendix C Hyperparameter investigations to unveil the most significant factor of degradation
@@ C.1 Training set size
C.2 Number of images per prompt
@@
C.3 Mixing real images to synthetic sets
Many previous works suggest augmenting the training set with real images to mitigate degradation during iterative training. We investigated whether mixing images from the original training set into the synthetic set at each iteration could alleviate this issue. At each iteration, images are randomly replaced with corresponding real images. Figure 14 and 15 show how degradation in the Chain of Diffusion varies when $50$% and $90$% of images are replaced for Pokemon and CelebA-1k datasets, respectively. Notably, even $10$% synthetic images are sufficient to induce degradation, and $50$% replacement rarely slows it down. CelebA-1k dataset appears to be significantly more susceptible to degradation.
@@ C.3 Mixing real images to synthetic sets
Figure 15: Chain of Diffusion on CelebA-1k dataset with real images randomly replacing synthetic images at each iteration. A $50$% replacement rarely slows degradation, while $10$% synthetic images are sufficient to initiate it. It suffers from more severe degradation than Pokemon dataset as compared with Figure 14.
C.4 Prompt Set
@@ C.5 U-Net and Text-Encoder
C.6 Number of diffusion steps
@@ C.9 CLIP Skip
C.10 Adding Gaussian noise to the original training set
@@ C.11 Stable Diffusion XL
Figure 22: Chain of Diffusion of SDXL on Pokemon dataset. Due to insufficient investigation into optimal hyperparameters for finetuning SDXL, our experiments largely rely on those from Stable Diffusion v1.5.
Appendix D Quantitative Trait Modeling
D.1 Proof of Theorem 1
@@ D.2 Simulation setup
D.2.1 Computing the radial sum of power spectra.
@@
D.2.2 Simulation parameters matching different CFG scales.
@@
D.2.3 ReDiFine and mutations.
ReDiFine combines condition drop finetuning and CFG scheduling, inspired by the mutation mechanism that compensates for distribution shift and maintains genetic distributions in population genetics. We apply two modifications to our theoretical analysis of Section 4—adding mutation variance to heritability and smoothing truncations—to simulate the effects of ReDiFine in the Chain of Diffusion. Specifically, we add the mutation standard deviation $\sigma_{M}$ of $0.1$ to heritability as:
@@
Appendix E ReDiFine
@@
In addition to visual inspections for images generated using ReDiFine, we compare the quantitative results of ReDiFine to baselines with different CFG scales across FID and CLIP score. Furthermore, we compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the text prompt as the fidelity metric applicable for text-to-image generation,
${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (17)
to evaluate each iteration $k$. SFD overcomes the problem of FID being sensitive to the number of images to compare. DiNOv2 features are used to compute SFD. The results, shown in Figure 28, demonstrate that ReDiFine achieves performance comparable to the optimal CFG scales ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly) across different datasets and metrics.
@@
Appendix F Ablation study
@@
F.1 Condition drop finetuning
@@ Appendix G Analysis
G.1 Latent analysis
Figure $3$(a) illustrates how the distribution of latent values evolves across different iterations. The histograms show the final latent vectors before decoding into pixel space, comparing various CFG scales and ReDiFine. For a CFG of $1.0$, the latent distribution rapidly converges into a Gaussian-like shape, with its variance shrinking over iterations. This behavior is consistent with previous work [3, 16, 11], which theoretically predicted that the self-consuming loop progressively trims the tails of the distribution, reducing output diversity until it collapses to a single mode. We hypothesize that this narrowing in the latent space leads to blurrier, more homogeneous outputs in pixel space. Conversely, at a CFG scale of $7.5$, the latent distribution develops longer tails and tends toward a more uniform spread across space. A CFG scale of $2.5$, which demonstrates the best reusability among the three, better preserves the latent distribution over iterations. ReDiFine further enhances this preservation, maintaining the histogram from the first to the last iteration, thus achieving both high fidelity in the first iteration and better reusability.
G.2 Differences between conditional and unconditional scores
@@
G.3 Power spectra of 2D Fourier transforms
@@
G.4 Fingerprints for forensic analysis

### v3

@@ B.1 Hyperparameters
We finetune Stable Diffusion v1.5 [32] using LoRA [33] at each iteration, with ft-MSE [57] as a fixed VAE to project images into latent space. Horizontal flip is the only image augmentation applied. Our implementation follows [34] and is built on PyTorch v2.2.2 [58], with torchvision 0.17.2, running on CUDA 12.4 using NVIDIA A-100 and L40S GPUs. Each finetuning and generation takes about 2 hours for  1k images. All default hyperparameters are listed in Table 1.
@@ B.2 Datasets
Butterfly [37] includes $832$ real picture images across $8$ species, using BLIP captions (length $50$-$75$ words) and species descriptions.
B.3 Evaluation metrics
FID
The Fréchet Inception Distance (FID) [39] is a widely used metric for evaluating the quality of generative models. It measures the similarity between the distributions of generated images and real images by comparing their feature representations extracted from a pretrained Inception network. Following [22], we use DiNOv2 feature extractor [40] instead of Inception, as it aligns better with human evaluation for high-quality images.
CLIP score
The CLIP score is a metric used to evaluate the alignment between an image and a text description, leveraging the pretrained CLIP (Contrastive Language-Image Pre-training) model [42]. CLIP is designed to encode both images and text into a shared embedding space, where semantically related image-text pairs have high cosine similarity. Based on this embedding space, the CLIP score measures the text-alignment capability of text-to-image generative models by computing cosine similarities of generated images and text descriptions used for generations.
SFD
We suggest and compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the same text prompt as the fidelity metric applicable for text-to-image generation. It computes text-prompt-wise similarities of the real image and generated image, more directly comparing two pairs of images while other metrics compares the overall distribution of datasets. We calculate
${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (6)
to evaluate each iteration $k$. SFD overcomes the problem of FID being sensitive to the number of images to compare.
Recall
As a key metric for evaluating the diversity of generative models, recall measures the fraction of the real data distribution that is captured by the generated data distribution. Following the method proposed by [61], we utilize the DiNOv2 feature extractor and set the number of neighbors to 5 for our evaluations.
Appendix C Different CFG Scales
This section presents images from Chain of Diffusion at various CFG scales for the four datasets. Figure 6, 7, 8, and 9 corresponds to Pokemon, CelebA-1k, Kumapi, and Butterfly, respectively. Each dataset shows results for five CFG scales, covering high, medium, and low values. The optimal medium CFG scale is $2.5$ for Pokemon and Kumapi, and $1.5$ for CelebA-1k and Butterfly. Lower-than-optimal CFG scales lead to low-frequency degradations, while higher-than-optimal scales result in high-frequency degradations with saturated colors and repetitive patterns. Notably, the optimal CFG for Pokemon and Kumapi causes severe degradation in CelebA-1k and Butterfly.
@@ Appendix C Different CFG Scales
Appendix D Hyperparameter investigations to unveil the most significant factor of degradation
@@ D.1 Training set size
D.2 Number of images per prompt
@@
D.3 Mixing real images to synthetic sets
Many previous works suggest augmenting the training set with real images to mitigate degradation during iterative training. We investigated whether mixing images from the original training set into the synthetic set at each iteration could alleviate this issue. At each iteration, images are randomly replaced with corresponding real images. Figure 12 and 13 show how degradation in the Chain of Diffusion varies when $50$%, $90$%, $95$%, $98$% and $99$% of images are replaced for Pokemon and CelebA-1k datasets, respectively. Notably, even $5$% synthetic images are sufficient to induce degradation, and $50$% replacement rarely slows it down. CelebA-1k dataset appears to be significantly more susceptible to degradation. Importantly, using only $5\%$ and $2\%$ synthetic images is sufficient to trigger model collapse in Pokemon and CelebA-1k datasets. This demonstrates the catastrophic impact of synthetic data on diffusion finetuning and raises the question of whether synthetic data detection can be a safe solution.
@@ D.3 Mixing real images to synthetic sets
Figure 13: Chain of Diffusion on CelebA-1k dataset with real images randomly replacing synthetic images at each iteration. A $50$% replacement rarely slows degradation, while $5$% synthetic images are sufficient to initiate it. It suffers from more severe degradation than Pokemon dataset as compared with Figure 12.
D.4 Prompt Set
@@ D.5 U-Net and Text-Encoder
D.6 Number of diffusion steps
@@ D.9 CLIP Skip
D.10 Adding Gaussian noise to the original training set
@@ D.11 Stable Diffusion XL
Figure 20: Chain of Diffusion of SDXL on four datasets. Due to insufficient investigation into optimal hyperparameters for finetuning SDXL, our experiments largely rely on those from Stable Diffusion v1.5.
D.12 Data Accumulation
Data accumulation experiments aim to investigate whether concept overfitting and disappearing are major reasons for model collapse. The training set for iteration $t$ is the combination of all previously generated sets, including the original training set. The number of training epochs is controlled accordingly to maintain the total number of updates.
Figure 21: Chain of Diffusion on Pokemon dataset when training set is accumulated from previous iterations. All concepts from previous iterations are preserved for finetuning.
Figure 22: Chain of Diffusion on CelebA-1k dataset when training set is accumulated from previous iterations. All concepts from previous iterations are preserved for finetuning.
Appendix E Quantitative Trait Modeling
E.1 Wright-Fisher model - Genetic drift and fixation
The Wright-Fisher model describes the evolution of discrete allele frequencies in a finite population. We assume distinguishable iterations (no overlapping individuals) and no selection or mutation. In each iteration, all individuals from the previous iteration (the parents) are replaced by new individuals (the offspring). The offspring are sampled from the parental population via random sampling with replacement; for simplicity, we assume a haploid model.
Consider a population of constant size $N$ and two alleles, $A$ and $a$. If allele $A$ is present in $i$ individuals in the parent interaction, the probability that it appears in $j$ individuals in the offspring iteration is given by the binomial distribution:
$P_{ij}=\binom{N}{j}({i\over N})^{j}(1-{i\over N})^{N-j},0\leq i,j\leq N.$ (7)
This process defines a Markov chain, where the state in each iteration depends only on the state in the previous one. The Wright-Fisher model thus characterizes the stochastic dynamics of allele frequencies across iterations.
Certain states in a Markov chain are absorbing—that is, once entered, they cannot be exited. In the Wright-Fisher model, the fixation or loss of an allele corresponds to such absorbing states. Once the population becomes entirely composed of a single allele, the other allele cannot reappear under the assumptions of no mutation or migration. Importantly, in a finite population, the process is guaranteed to eventually reach one of these two absorbing states.The probability that allele $A$ becomes fixed depends solely on its initial frequency in the population.
E.2 Proof of Theorem 4.1
@@ E.3 Simulation setup
E.3.1 Computing the radial sum of power spectra.
@@
E.3.2 Simulation parameters matching different CFG scales.
@@
E.3.3 ReDiFine and mutations.
ReDiFine combines condition drop finetuning and CFG scheduling, drawing inspiration from genetic mutation mechanisms that counter distributional shifts and help maintain genetic diversity in population genetics. In the Chain of Diffusion, training and generation can be viewed analogously to reproduction and natural survival. ReDiFine modifies both stages to preserve image quality across iterations.
The two core components—condition drop finetuning and CFG scheduling—are rooted in biological analogies. Condition drop finetuning, akin to mutation, introduces controlled variability to ensure that the offspring distribution (the next iteration) remains aligned with the parent distribution (the previous iteration). This mitigates over-specialization (generated images becoming too uniform or isolated) and prevents degeneration under strong selection pressures. CFG scheduling, inspired by natural selection, dynamically adjusts the CFG scale during generation, promoting diversity by allowing broader phenotypic expression while retaining directed evolution toward desired traits. By integrating these biologically inspired strategies, ReDiFine improves the stability of the Chain of Diffusion and prevents catastrophic degradation across iterations.
We apply two modifications to our theoretical analysis of Section 4—adding mutation variance to heritability and smoothing truncations—to simulate the effects of ReDiFine in the Chain of Diffusion. Specifically, we add the mutation standard deviation $\sigma_{M}$ of $0.1$ to heritability as:
@@
Appendix F ReDiFine
F.1 Visual Inspections
@@
F.2 Further iterations
We conduct additional experiments to compare the baseline with the optimal CFG scale and ReDiFine over extended iterations. As shown in Figure 28, ReDiFine consistently generates images of similar quality up to 12 iterations, whereas the optimally tuned CFG scale fails to sustain image quality. This decline suggests that repeated hyperparameter searches are necessary to identify suitable CFG scales for subsequent iterations. Such an approach becomes increasingly impractical as the number of iterations grows, highlighting the limitations of relying on the optimal CFG scale to mitigate model collapse.
Figure 28: Chain of Diffusion of baseline with the optimal CFG scale and ReDiFine for more iterations. The generation quality of ReDiFine is preserved for additional iterations while optimally found CFG scale $2.5$ fails to maintain the image qualities. Similar high-frequency degradation is observed.
F.3 Cross-domain Data
We additionally conduct experiments when the training dataset is the cross-domain set of four datasets (Pokemon, CelebA-1k, Kumapi, and Butterfly) to investigate whether having a broader range of concepts impact model collapse. The merged dataset serves as the original training set, while the combined captions are used for image generation at each iteration. The results, presented in Figure 29, show that despite the increased number of images and the inclusion of diverse concepts and domains, model collapse persists at both low and high CFG scales. Moreover, no single CFG scale (e.g., $1.5$ or $2.5$) can consistently produce high-quality, reusable images across all datasets, highlighting the limitations of relying on an optimal CFG scale for diverse domains. In contrast, ReDiFine leverages the increased conceptual diversity in the original training set, generating more reliable images across all datasets. To ensure a fair comparison, we control the number of training epochs to maintain consistent updates across experiments.
Figure 29: Chain of Diffusion with the multi-domain set of datasets (Pokemon, CelebA-1k, Kumapi, and Butterfly) used for finetuning. Model collapse remains evident at low and high CFG scales. A single CFG scale (e.g., $1.5$ or $2.5$) fails to achieve optimal performance across both the Pokemon and CelebA-1k datasets. In contrast, ReDiFine successfully generates high-quality, reusable images simultaneously. While images for Kumapi and Butterfly datasets are not displayed due to space constraints, they are included in the finetuning process along with the other datasets.
F.4 Iterative Retraining
Some prior works on model degeneration examine scenarios in which a single model is continually trained on synthetic data it has generated. To adapt our Chain of Diffusion framework to this setting, we consider a setup where the same model is finetuned iteratively across multiple iterations. At each iteration, the model generates a fixed number of images using a predefined prompt set, and these generated images are then used to further finetune the model. Figure 30 demonstrates how images degrade under this setting across different CFG scales and ReDiFine. While severe degradation is observed for low and high CFG scales, the optimal CFG scale and ReDiFine are able to mitigate model collapse, generating high-quality images. This indicates that the effect of ReDiFine is maintained even when a single model is continually finetuned.
Figure 30: Chain of Diffusion where a single model is continually finetuned across multiple iterations. Model collapse is observed consistently at CFG scales of 1.0 and 7.5, while the baseline with a CFG scale of 2.5 and ReDiFine effectively mitigate model collapse. This demonstrates that model collapse is a universal phenomenon across different settings, and the effect of ReDiFine is effective robustly.
F.5 Quantitative results
In addition to visual inspections for images generated using ReDiFine, we compare the quantitative results of ReDiFine to baselines with different CFG scales across FID, CLIP score, SFD, and recall. Details about each metric are provided in Appendix B.3. DiNOv2 features are used to compute FID, recall, and SFD. We follow [61] to compute recall and set the number of neighbors for computing recall $5$. The results, shown in Figure 31, demonstrate that ReDiFine achieves performance comparable to the optimal CFG scales ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly) across different datasets and metrics.
@@ F.5 Quantitative results
(m) Pokemon, recall.
(n) CelebA-1k, recall.
(o) Kumapi, recall.
(p) Butterfly, recall.
@@
Appendix G Ablation study
@@
G.1 Condition drop finetuning
@@ Appendix H Analysis
H.1 Latent analysis
Figure 36(a) illustrates how the distribution of latent values evolves across different iterations. The histograms show the final latent vectors before decoding into pixel space, comparing various CFG scales and ReDiFine. For a CFG of $1.0$, the latent distribution rapidly converges into a Gaussian-like shape, with its variance shrinking over iterations. This behavior is consistent with previous work [3, 16, 11], which theoretically predicted that the self-consuming loop progressively trims the tails of the distribution, reducing output diversity until it collapses to a single mode. We hypothesize that this narrowing in the latent space leads to blurrier, more homogeneous outputs in pixel space. Conversely, at a CFG scale of $7.5$, the latent distribution develops longer tails and tends toward a more uniform spread across space. A CFG scale of $2.5$, which demonstrates the best reusability among the three, better preserves the latent distribution over iterations. ReDiFine further enhances this preservation, maintaining the histogram from the first to the last iteration, thus achieving both high fidelity in the first iteration and better reusability.
H.2 Differences between conditional and unconditional scores
@@
H.3 Power spectra of 2D Fourier transforms
@@
H.4 Fingerprints for forensic analysis